Here's a problem that I can not figure out.
8-x/3=6(x+1)/7
I really appreciate you helping me out.

Answer by checkley71(8403)

8-X/3=6(X-1)/7 OR (24-X)/3=(6X-6)/7 OR CROSS MULTIPLYING WE GET 3(6X-6)=7(24-X)
18X-18=168-7X 18X+7X=168+18 OR 25X=186 OR X=186/25 X=7.44
8-7.44/3=6(7.44-1)/7 OR (24-7.44)/3=6(7.44-1)/7 16.56/3=(6*6.44)/7 OR
5.52=38.64/7 OR 5.54=5.54